OpenStudy (anonymous):

7y - 3y - 14 = 14 y =

OpenStudy (muzzack):

Rearrange : Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : 7*y-3*y-14-(14)=0 Step by step solution : Step 1 : Pulling out like terms : 1.1 Pull out like factors : 4y - 28 = 4 • (y - 7) Equation at the end of step 1 : 4 • (y - 7) = 0 Step 2 : Solve 4•(y-7) = 0 Equations which are never true : 2.1 Solve : 4 = 0 This equation has no solution. A a non-zero constant never equals zero. Solving a Single Variable Equation : 2.2 Solve : y-7 = 0 Add 7 to both sides of the equation : y = 7 One solution was found : y = 7